# Convert Language Model into a Value-based Strategic Planner

### v4

@@ Abstract
Emotional support conversation (ESC) aims to alleviate the emotional distress of individuals through effective conversations. Although large language models (LLMs) have obtained remarkable progress on ESC, most of these studies might not define the diagram from the state model perspective, therefore providing a suboptimal solution for long-term satisfaction. To address such an issue, we leverage the Q-learning on LLMs, and propose a framework called straQ*. Our framework allows a plug-and-play LLM to bootstrap the planning during ESC, determine the optimal strategy based on long-term returns, and finally guide the LLM to response. Substantial experiments on ESC datasets suggest that straQ* outperforms many baselines, including direct inference, self-refine, chain of thought, finetuning, and finite state machines.
@@ 3.1 Task Definition
The problem of emotional-support conversation (ESC) can be characterized by an interleaved sequence of seeker $query$ and supporter $response$. To strengthen emotional-support performance, recent studies Liu et al. (2021); Rashkin et al. (2019) enhance the data content by augmenting the set of support strategies $\mathcal{A}$ and seeker emotions $\mathcal{E}$. For each conversation session, the background $description$ is also annotated on the session-level. Such augmented ESC can then be described as
@@
in which $desc$ and $resp$ are the abbreviations of $description$ and $response$, and $T$ is the total number of conversation turns. At turn $t$, we denote the conversation history as
@@ 3.3 Implementation on Language Models
[…]
@@ 3.4 Reward Definitions
(2) Imitation: we consider each ($s$, $a$) pair from the dataset is an expert demonstration, therefore, always assigned with $r$ of $+1$. To amplify the distribution, we randomly sample a different $a$ and assign with $r$ of $-1$. The positive-negative ratio is 1:1. Since this manner imitates the positive samples directly, this variant is called straQ*-imit.
@@ 4.1 Setting
Llama3.2-1B-instruct AI@Meta (2024) is employed as the base model. Training is conducted on OpenRLHF Hu et al. (2024), with the learning rate of $65.0e-$ 6, window length of 2048, batch size of 64, and epoch of 4. The target network update frequency is set to 10, the replay buffer size is 12,000, and $\gamma=0.85$.
@@ 4.2 Evaluation Methods
Methods Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$
@@ 4.4 Discussions
Method Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$
@@
supporter Res.& Par. (I) So as far as I understand that you had issues with your wife due to your drinking. Am I right? 3 7.6
Direct Res.& Par. (I) So, it sounds like there’s been a bit of tension in your marriage lately due to your drinking habits, and it even led to a confrontation that’s caused a rift between you and your wife. Can you tell me more about what’s been going on with your drinking? 4 8.0
straQ*-distill Res.& Par. (I) It sounds like your drinking habits have been causing difficulties in your marriage for a while. 3 10.1
@@
supporter Aff.& Rea. (II) That is really a serious problem. I know it must be very tough time for you. Hope you are trying your best to save your marriage. 2 5.4
Direct Res.& Par. (I) Just to make sure I understand correctly, it sounds like there was a confrontation with your wife about your drinking habits, and now you two aren’t speaking. Is that a fair summary of what happened? 3 4.7
straQ*-distill Ref. (II) That must be really difficult for you. I can imagine how much stress and worry this situation has brought into your life. 4 8.3
@@
supporter Ref. (II) That’s is a great initiative. I know it is tough for you but not impossible. I can feel how hard you are trying t save your relationship. I appreciate you for that. 4 4.0
Direct Res.& Par. (I) Just to make sure I understand correctly, you’ve been trying to cut down on your drinking to address your wife’s concerns, but despite your efforts, the issue still seems to be causing tension in your marriage, and now you’re not talking to each other. Is that a correct summary? 2 2.0
straQ*-distill Aff.& Rea. (III) That’s a great step forward! It takes a lot of strength and commitment to make such a change, and you should be proud of yourself. 5 5.0
@@
Strategy Acc $\uparrow$ $\mathcal{Q}$uparrow $\mathcal{B}$downarrow B-2 $\uparrow$ R-L $\uparrow$ Dist-2 CIDEr
SFT Que. 57.52 48.24 1.60 9.37 21.88 64.42 34.22
@@ Appendix A Dataset details
Topic I hate my job but I am scared to quit and seek a new career.
Query {history} seeker: Seriously! What I’m scare of now is how to secure another job.
@@
Strategy Reflection of feelings
Response supporter: I can feel your pain just by chatting with you.
@@ Appendix B Detailed Prompts
Emotion: {$e$}
Description: {$desc$}
@@
Emotion: {$e$}
Description: {$desc$}
@@
Emotion: {$e$}
Description: {$desc$}

### v5

@@ Abstract
Emotional support conversation (ESC) aims to alleviate the emotional distress of individuals through effective conversations. Although large language models (LLMs) have obtained remarkable progress on ESC, most of these studies might not define the diagram from the state model perspective, therefore providing a suboptimal solution for long-term satisfaction. To address such an issue, we leverage the Q-learning on LLMs, and propose a framework called straQ*. Our framework allows a plug-and-play LLM to bootstrap the planning during ESC, determine the optimal strategy based on long-term returns, and finally guide the LLM to response. Substantial experiments on ESC datasets suggest that straQ* outperforms many baselines, including direct inference, self-refine, chain of thought, finetuning, and finite state machines. Our implementation is available at GitHub.
@@ 3.1 Task Definition
The problem of emotional-support conversation (ESC) can be characterized by an interleaved sequence of seeker queryquery and supporter responseresponse. To strengthen emotional-support performance, recent studies Liu et al. (2021); Rashkin et al. (2019) enhance the data content by augmenting the set of support strategies $\mathcal{A}$ and seeker emotions $\mathcal{E}$. For each conversation session, the background descriptiondescription is also annotated on the session-level. Such augmented ESC can then be described as
@@
in which descdesc and respresp are the abbreviations of descriptiondescription and responseresponse, and $T$ is the total number of conversation turns. At turn $t$, we denote the conversation history as
@@ 3.3 Implementation on Language Models
Prompt Template Description: {descdesc}    User’s emotion: {ee} History: {hh}    Query: {queryquery} Please select the best strategy: (1) {strategy1strategy1}  (2) $\cdots$  (K) {strategyKstrategyK}
@@ 3.4 Reward Definitions
(2) Imitation: we consider each ($s$, $a$) pair from the dataset is an expert demonstration, therefore, always assigned with $r$ of +1+1. To amplify the distribution, we randomly sample a different $a$ and assign with $r$ of 1-1. The positive-negative ratio is 1:1. Since this manner imitates the positive samples directly, this variant is called straQ*-imit.
@@ 4.1 Setting
Llama3.2-1B-instruct AI@Meta (2024) is employed as the base model. Training is conducted on OpenRLHF Hu et al. (2024), with the learning rate of 5.0e65.0e-6, window length of 2048, batch size of 64, and epoch of 4. The target network update frequency is set to 10, the replay buffer size is 12,000, and $\gamma=0.85$.
@@ 4.2 Evaluation Methods
Methods Acc $\uparrow$ $\mathcal{Q}$ $\uparrow$ $\mathcal{B}$ $\downarrow$ B-2 $\uparrow$ R-L $\uparrow$
@@ 4.4 Discussions
Method Acc $\uparrow$ $\mathcal{Q}$ $\uparrow$ $\mathcal{B}$ $\downarrow$ B-2 $\uparrow$ R-L $\uparrow$
@@
supporter
Res.& Par.
So as far as I understand that you had issues with your wife due to your drinking. Am I right? 3 7.6
Direct
Res.& Par.
So, it sounds like there’s been a bit of tension in your marriage lately due to your drinking habits, and it even led to a confrontation that’s caused a rift between you and your wife. Can you tell me more about what’s been going on with your drinking? 4 8.0
straQ*-distill
Res.& Par.
It sounds like your drinking habits have been causing difficulties in your marriage for a while. 3 10.1
@@
supporter
Aff.& Rea.
That is really a serious problem. I know it must be very tough time for you. Hope you are trying your best to save your marriage. 2 5.4
Direct
Res.& Par.
Just to make sure I understand correctly, it sounds like there was a confrontation with your wife about your drinking habits, and now you two aren’t speaking. Is that a fair summary of what happened? 3 4.7
straQ*-distill
Ref.
That must be really difficult for you. I can imagine how much stress and worry this situation has brought into your life. 4 8.3
@@
supporter
Ref.
That’s is a great initiative. I know it is tough for you but not impossible. I can feel how hard you are trying t save your relationship. I appreciate you for that. 4 4.0
Direct
Res.& Par.
Just to make sure I understand correctly, you’ve been trying to cut down on your drinking to address your wife’s concerns, but despite your efforts, the issue still seems to be causing tension in your marriage, and now you’re not talking to each other. Is that a correct summary? 2 2.0
straQ*-distill
Aff.& Rea.
That’s a great step forward! It takes a lot of strength and commitment to make such a change, and you should be proud of yourself. 5 5.0
@@
Strategy Acc $\uparrow$ $\mathcal{Q}$ $\uparrow$ $\mathcal{B}$ $\downarrow$ B-2 $\uparrow$ R-L $\uparrow$ Dist-2 CIDEr
SFT
Que. 57.52 48.24 1.60 9.37 21.88 64.42 34.22
@@ Appendix A Dataset details
Topic
I hate my job but I am scared to quit and seek a new career.
Query
{history}
seeker: Seriously!
What I’m scare of now is how to secure another job.
@@
Strategy
Reflection of feelings
Response
supporter: I can feel your pain just by chatting with you.
@@ Appendix B Detailed Prompts
Emotion: {ee}
Description: {descdesc}
@@
Emotion: {ee}
Description: {descdesc}
@@
Emotion: {ee}
Description: {descdesc}